sailboat sails 2.0 km east, then 4.0 km southeast, then an additional distance in an unknown direction. Its final position is 6.0 km directly east of the starting point. Find the magnitude and direction of the third leg of the journey graphically. (Scale:


1
Expert's answer
2021-03-31T07:13:55-0400

The direction is shown in figure (x).



According to the cosine law:


"x = \sqrt{4^2 + 4^2 - 2\cdot4\cdot 4\cos 45\degree } \approx 3.06 km"

Answer. 3.06 km.